OpenStudy (anonymous):

Select the two values of x that are roots of the given polynomial below. 3x2 + 11x - 4

OpenStudy (anonymous):

3x^2+12x-x-4=0 3x(x+4)-(x+4)=0 (3x-1)((x+4)=0 X=1/3 or x= -4